OpenStudy (anonymous):

Josh believes the Spanish club students at his school have an unfair advantage in being assigned to the Spanish class they request. He asked 500 students at his school the following questions: "Are you in the Spanish club?" and "Did you get the Spanish class you requested?" The results are shown in the table below: Spanish Club Not in Spanish Club Total Received Spanish class requested 265, 100, 365 Did not get Spanish class requested70, 65, 135 Total 335, 165, 500 Help Josh determine if all students at his school have an equal opportunity to get the Spanish class they requested.

OpenStudy (anonymous):

Spanish Club Not in Spanish Club Total Received Spanish class requested 265 100 365 Did not get Spanish class requested 70 65 135 Total 335 165 500

OpenStudy (usman1995):

i don't get ur question

OpenStudy (anonymous):

Based on the above data, does everyone have an equal chance of getting into the Spanish class that they want?

OpenStudy (anonymous):

P(students in spanish club who recieved spanish class) = 265/365 = 53/73 P(students NOT in spanish club who recieved spanish class) = 100/365 = 20/73 so all students dont have equal opportunity to get the Spanish class they requested

OpenStudy (usman1995):

oh sorry

OpenStudy (usman1995):

oh i havnt got the idea of question havent read it all was just looking column wise lol

OpenStudy (anonymous):

@Usman1995 lagta tu bhool gea probability

OpenStudy (usman1995):

oye nai mene swaaal sara nai para

OpenStudy (anonymous):

@CO_oLBoY 1. Thank-you so much 2. Is there any way that you can explain that in a little more depth? If, no then, ok.

OpenStudy (anonymous):

ok @GreenEyedGirl i'll explain

OpenStudy (usman1995):

265 students received spanish class request and they were in the club too and total no. of students who received request is 365 so there probability is 265/365

OpenStudy (usman1995):

whereas students who received request for class but r not in spanish club are 100 since there are total no. of 365 students who got the request so the probability for the students not in spanish club will be 100/365

OpenStudy (usman1995):

based of the answers of @shami that's the best explanation

OpenStudy (anonymous):

Thank-you all soooooooo much.

OpenStudy (anonymous):

the question is saying that we have to find that all the students of Josh's class wether they r in club or not can attend spanish class or not. So we just take the data of people who can attend the spanish class and take out the probilities of that. for finding the probabilty we divide the no. of students of each group who can attend the class by the total no. of students who atteded..

OpenStudy (anonymous):

if both group probabilities are same then all students have equal opportunity to get the Spanish class otherwise they dont have equal opportunity

OpenStudy (usman1995):

NP

OpenStudy (anonymous):

@GreenEyedGirl got it?

OpenStudy (anonymous):

If I open another question, I can give you another medal. Are you willing to help?

OpenStudy (anonymous):

@Usman1995 my username is CO_oLBoY not shami

OpenStudy (anonymous):

@GreenEyedGirl i'll be delighted to help

OpenStudy (usman1995):

acha g kar tinnu koi roti nai ponda ty wadda tu cool ,, apni asliyyat buri lagti hai kia tujy

OpenStudy (anonymous):

o ni me keh ra tha k tune kisi or ko tag kr dya

OpenStudy (anonymous):

tag krne k leye username hi likha kr @Usman1995

OpenStudy (usman1995):

lol